The resistance of a metal is given by R = V/1, , where V is potential difference and 1 is the current. In a circuit the potential difference across resistance is V = (8 + 0.5) V and current in resistance, I = (2 + 0.2)A. What is the value of resistance with its percentage error?


        (a)    4W + 16.25%


        (b)    (4 + 0.7)W


        (c)    4W + 0.7%


        (d)    4W + 7%



Grade:10

1 Answers

Navjyot Kalra
10 years ago

(a)

    Actual R = 8/2 = 4

    % error in R = % error in V + % erro in 1

    = 0.5/8×100%+0.2/2×100%=16.25%

    R = 4W + 16.25%